Turkish grammar

David Pierce

2005.05.02

Contents

0 Preliminaries 1 0.0 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 0.1 Alphabet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 0.2 Sounds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 0.3 Writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1 Parts of speech 5 1.0 Words . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.1 Nouns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.2 Adjectives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.3 Numerals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.4 Pronouns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.5 Adverbs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.6 Particles, conjunctions and postpositions . . . . . . . . . . . . 11 1.7 Verbs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

1.7.1 Stems . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 1.7.2 Bases . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2 Glossary 18

0 Preliminaries

0.0 Introduction

As a student of Turkish, I make these notes in an effort to understand the logic of the Turkish language. This is not the account of an expert; as should be quite clear, this is not anything like a complete exposition of Turkish grammar. The information here is mostly from secondary sources: I mainly use Lewis’s Turkish Grammar [1], but have also consulted [2]. I have made some use of primary sources, namely, Turkish speech and writing as they appear in the course of my life in Turkey.

I shall propose some technical terms that are different from Lewis’s. Since my account is written, it will use the Turkish alphabet; therefore

my account illustrate features like vowel-harmony that are reflected in spelling.

0.1 Alphabet

Like English, modern Turkish (since 1928) uses an alphabet derived from the 23-letter Latin alphabet

A, B, C, D, E, F, G, H, I, K, L, M, N, O, P, Q, R, S, T, V, X, Y, Z.

English gets 3 more letters by introducing:

• the variant J of I, and

• the variants U and W (double-U) of V.

The letter-forms given above are called upper case, to distinguish them from the lower case (a, b, c and so on).

The 29-letter Turkish alphabet can be derived from the 26-letter English one by:

• eliminating Q, X and W;

• deriving from C, G, O, S and U the letters Ç, Ğ, Ö, Ş and Ü respectively, by adjoining cedilla, breve or umlaut;

• replacing I with two letters, I and İ (distinguished in both upper and lower case by whether a dot is present: the lower-case forms are therefore ı and i respectively).

The letter Ğ is called yumuşak ge (soft G), and it never begins a word. The letters that are ‘new’ with respect to the English alphabet can be understood as derived from the old letters by means of diacritical marks. However, these marks become integral parts of the new letters. It is not the shape of the mark, but its presence at the top or bottom of the letter that is used to distinguish the letter from its ‘twin’. For example, one Turkish dairy styles itself Süta\begin{matrix}s \\ . \end{matrix} (they use a dot instead of a cedilla under the s), and there is a bus company called Uludag̀ (taking their name from the mountain at Bursa, the first capital of the Ottoman Empire; they use a grave accent rather than a breve over the g).

In the alphabetical order, I precedes İ, and the other new letters follow their ‘twins’. Hence the Turkish alphabet is

A, B, C, Ç, D, E, F, G, Ğ, H, I, İ, J, K, L,

M, N, O, Ö, P, R, S, Ş, T, U, Ü, V, Y, Z.

Because of the distinction between the letters ı and i in Turkish, neither of these should be joined in a ligature to a preceding f. In English, the other word for a movie is film: the word is usually printed thus, with three characters, namely fi, i and m. But film is also a Turkish word, and as such is printed with four distinct letters. (Also, for example, fil means elephant or chess-bishop, while fındık means hazelnut.)

In Turkey, I have seen a capital J written with a dot. But j is rare in Turkish anyway, being used only in foreign words.

0.2 Sounds

The name of a vowel is itself, and the name of a consonant (besides Ğ, yumuşak G) is itself plus e.

The number of vowels is eight, that is, 2^{3}, since each vowel is determined by the quality it possesses from each of three pairs: thick/thin, flat/round and narrow/wide. These are literal translations of the usual Turkish terms

(kalın/ince, düz/yuvarlak, dar/geniş), but commonly in English the

terms are:

• back/front,

• unround/round,

• close/open.

The vowels can then be tabulated:

open  close
back front back front
unround a  e  ı  i
round  o  ö  u  ü

The sound of a can be spelled in English by uh; close the mouth more to get ı. The ö and ü are as in German; their sounds are found in French as well.

Further distinctions are possible, mainly because Turkish retains borrowings from Arabic and Persian. A difference from the norm might be shown by a circumflex: so kar means snow, but k^ar, sounding something like kyahr, means profit.

The other 21 letters are consonants. Besides the ‘semi-vowel’ y, the consonants might be tabulated:

b/p v/f  m
d/t j/ş c/ç n l

z/s

g/k ğ/h  r

The rows here are intended to correspond to position of the lips and the tip of the tongue; the columns, to the flow of breath used to pronounce the consonant. (This particular table is my own invention though, and is not the result of careful phonological study.) The members of the pairs ∗/∗ are voiced and unvoiced respectively, or soft and strong in Turkish (yumuşak and sert). The main effect of ğ is to lengthen the preceding vowel. (So it functions like gh in English in changing fit into fight, although the English long I is not the same as the Turkish long İ.) The letter j is pronounced as in French. The sound of c is spelled in English by j or dge; the sounds of ç and ş, by ch and sh.

0.3 Writing

As in English, written Turkish is divided into sentences, and these into words. Some words are enclitic, being unaccented, but causing the previous syllable to be accented. (Syllabic accents are not very strong in Turkish though.) The vowel in an enclitic may also change according to the preceding vowel. This same vowel harmony is shown by many suffixes.

I shall indicate the variability of vowels with the following symbols:

• @ for an open unround vowel (a or e);

• # for a close vowel (ı, i, u or ü).

I have not seen such symbols used elsewhere in this way. (Perhaps using æ instead of @ would be more logical, but its appearance seems more confusing.) Used in a complete word (or sentence), each symbol resolves to a vowel that agrees as far as possible with the preceding vowel. In particular, after @, the only possibilities for # are the unround vowels ı and i.

The variability in consonants will not be indicated. However, when it begins a suffix appended to a word ending in an unvoiced (that is, strong) consonant, the letter d is unvoiced (‘strengthened’) to t. Also, terminal k changes to ğ when a suffix beginning with a vowel is added. (These changes affect the suffix -d#r and the termination -k mentioned below. Other such changes can occur, as between c and ç.)

Example. The question Avrupa + l# + l@ş + d#r + @m@ + d#k + l@r
+ #m#z + d@n m# + s#n#z? resolves to Avrupalılaştıramadıklarımız-

dan mısınız? meaning Are you one of those whom we could not Europeanize?

The symbol # may resolve to zero (nothing) when it begins a suffix placed after a vowel; alternatively, a consonant may be interposed, n, s or y, depending on the situation. Sometimes the y narrows the preceding @ to #.

1 Parts of speech

1.0 Words

Not every word appears in a dictionary, even if the word is well-formed; the word may be inflected (or otherwise derived) from a dictionary-word according to standard rules. Otherwise, we might say that the same word may be used in many forms, only one or a few of which appear in the dictionary. For example, in English, we can say that man and men are two words, or else that they are two forms of the same word. Inflected forms are few in English; in Turkish they are many.

Words in a Turkish dictionary can be assigned to some of the same parts of speech as English words: noun, pronoun, adjective, numeral, verb, adverb and conjunction. Instead of prepositions, but fulfilling some of the same functions, Turkish has postpositions. Some grammatical functions are served by particles.

Some Turkish words used in examples like the following are defined in a glossary below (in § 2):

Example. The postposition gibi corresponds to the preposition like. So, buz gibi means like ice, that is, ice cold. (It is a slogan on signs in Turkey advertising a certain American soft drink that is marketed all over the world.)

This reversal of order in passing from the one language to the other occurs in other ways. A Turkish word (as in the earlier example) may have parts which, in English, would appear as separate words in the opposite order. As in English though, adjectives generally precede the nouns they modify.

1.1 Nouns

Turkish nouns have no gender. To a noun in the dictionary can be added four (kinds of) suffixes, in the following order:

1. The plural suffix, -l@r.

2. A possessive suffix, indicating the number (singular or plural) of a possessor as well as its person (first, second or third):

1st  2d  3d
singular -(#)m  -(#)n  -(s)#
plural -(#)m#z -(#)n#z -(l@r)#

3. A case-ending, roughly as in Latin:

• definite-accusative: -(y)#;

• genitive: -(n)#n;

• dative: -(y)@;

• locative: -d@;

• ablative: -d@n.

4. A predicative suffix, asserting the distinctions of number and person that apply to the noun itself:

1st  2d  3d
singular -(y)#m -s#n  -(d#r)
plural -(y)#z -s#n#z -(d#r)(l@r)

Lewis refers to the possessive suffixes as personal. However, as will be seen, Turkish has five series of suffixes that indicate the attribute called ‘person’. For the particular series of suffixes in question here, I have chosen a name that gives more indication of their function. (We might also refer to these endings as ‘possessive personal suffixes’.) Note however that these endings do not make a noun ‘possessive’ in the English sense in which man’s, my and mine are possessive; they make a noun ‘possessed’. (More on this point is just below.)

Lewis refers to the predicative suffixes as type I, since there will also be types II, III and IV. I chose the name I did, since these endings turn nouns into predicates (and even complete sentences).

In the endings given above, parenthetical single letters are used or not, depending on which is more euphonious. (The parenthetical s, y and n are used only between vowels; the parenthetical # is used only between consonants.) The plural suffix is parenthetical above because it is generally not repeated:

Example. The word evleri can be analysed as evler-i his/her/its houses or as ev-leri their house; it may also mean their houses. (It can also be the definite-accusative of evler houses.)

The suffix -d#r is said to come, as early as the 11th century, from the aorist verb turur stands. Its use is imprecisely summarized this way: It can be left out if ambiguity will not result, and it can be used for emphasis (even in persons besides the third).

Example. Atlar  They are horses Attırlar They are surely horses Atlardır  They are the horses

Note then that a single sentence can be built up out of a single word:

Example. eviniz  your house evinizde  at your house Evinizdeyim I am at your house If one considers ev as syn-

onymous with odalar rooms, then one can contrive Odalarınızdayım I am in your rooms, with all four kinds of suffix.

The second-person plural can be used (as in French) for a polite singular. A noun without the plural suffix is not really singular, but numberless.

A noun without a case-ending (or, a noun in the absolute case) can be an indefinite direct object:

Example. Şiir okur  (S)he reads poetry Şiirler okur  (S)he reads poems Şiirleri okur (S)he reads the poems

If a numeral (§ 1.3) is present, then the plural suffix is (normally) not used:

Example. Beş baş koyun five head of sheep.

If one noun is genitive and is to possess another, then that other noun takes a possessive ending. The converse need not be the case, for relations that are not strictly possessive:

Example. Türkiye'nin cumhurbaşkanı president of Turkey Türkiye Cumhuriyeti  Republic of Turkey

In general though, when two nouns form a unit, then the second one has a possessive ending; the first can be in the absolute or the genitive case. Except for the lack of a possessive suffix in the Turkish sense, English is the same way: Elm Street, my friend’s street. (The last example but one is exceptional. One also has altın küpe gold earring; but here altın, as gold, can be understood as an adjective, so küpe need not become küpesi.)

Some suffixes make new nouns from old. Common are:

• -c#, indicating -monger, -dealer, -seller, or more generally a person involved with the thing;

• -l#k, making some kind of abstraction.

Example. sütçü milkman günlük diary işçi worker gecelik nightgown

1.2 Adjectives

Adjectives can behave as nouns, taking all of the same endings that nouns do. Nonetheless, they can generally be distinguished from nouns in the manner just alluded to: When they modify a noun, they do not force the noun to take a possessive suffix.

As noted earlier, adjectives do generally come before the nouns they modify. Some adjectives have intensive forms, which represent the only use of prefixes in Turkish:

Example. bembeyaz whiter than white; bomboş utterly void, çırçıplak stark naked.

Unlike some English adjectives, Turkish adjectives do not have comparatives and superlatives; rather, adverbs are used, as for English adjectives of more than one syllable:

daha more en  most
az  less en az least

Also, in a comparison, the noun to which the comparison is being done is in the ablative case (and then daha need not be used).

The suffix -l# makes a noun into an adjective indicating possession of or connexion with the noun.

Example. şekerli  sweet hızlı  swift Ankaralı ‘Ankarian’

Two important words can be called adjectives: var and yok, meaning existent and non-existent. Putting one of them after a noun is like putting there is (no) before a noun in English: Tanrı yok(tur) means There is no god. Also, they are used to assert or deny possession, since Turkish has no verb have: So, Atım var means My horse exists, that is, I have a horse.

According to Lewis, a sentence like Çocuğun babası yok The child has no father should be understood as having the subject çoğun and the predicate babası yok. (The reason lies in the kind of modifiers that babası can take here.)

1.3 Numerals

The cardinal numbers are named according to the conventional base-ten numeration system, with special names for single-digit multiples of ten and for certain powers of ten as in English (without exceptions like eleven and thirteen):

                             sıfır 0

  bir   iki    üç   dört  beş   altı    yedi   sekiz   dokuz

1  2  3  4  5  6  7  8  9

  on  yirmi  otuz  kırk  elli  altmış  yetmiş  seksen  doksan

10  20  30  40  50  60  70  80  90

                      yüz  bin   milyon

100 1000 1 000 000

Example. An astronomical unit, 149 597 870 kilometres, is yüz kırk dokuz

milyon beş yüz doksan yedi bin sekiz yüz yetmiş kilometre.

Ordinals are formed from cardinals by addition of -(#)nc#. There are also distributive numerals, formed by addition of (ş)@r:

Example. birinci first birer one each ikinci second ikişer two each

The word bir has some similarities to the English indefinite article a(n).

1.4 Pronouns

The personal pronouns may supplement the predicative suffixes mentioned above:

1st 2d 3d
singular ben sen o
plural biz siz onlar

Declension is as for nouns, with some irregularities:

• the dative of ben is bana, and of sen, sana;

• the genitive of ben is benim, and of biz, bizim;

• o is declined as if it were on.

The o is also a demonstrative pronoun, like that. The other demonstratives are bu this, and şu this or that (thing pointed to); these are declined as if they were bun and şun.

There are no relative pronouns in Turkish. (However, see below on ki. Also, as in English, the interrogative pronouns can function as relatives: the Turkish Kimsin bilmiyorum can be translated as I don’t know [the answer to] ‘Who are you?’, that is, I don’t know who you are.) The role of an English relative clause is taken in Turkish by participles. (The same can happen in English: the man who is walking can be the walking man; for the road where the man is walking, we can try the walked-on-by-the-man road, but it is unnatural.)

Interrogative pronouns are: kim who, ne what, hangi which, and kaç how many. The accusative and ablative of ne correspond to why; and ne zaman is when.

Added to the demonstratives and to ne, the suffix -r@ makes nouns of place, some of whose cases have corresponding English words, some obsolescent:

buraya hither oraya thither nereye whither?

       burada here orada there nerede where?
       buradan hence oradan thence nereden whence?

(Likewise we have şuraya, şurada, şuradan.)

1.5 Adverbs

Adjectives can function as adverbs, without addition of a suffix like the English -ly, but sometimes with reduplication.

1.6 Particles, conjunctions and postpositions

The enclitic particle m# turns a sentence into a question, withe emphasis on the preceeding word; but it does not follow a predicative ending; rather, this ending jumps to m#:

Example. Evde misin? Are you at home (and not somewhere else)?; Sen mi evdesin? Are you (and not somebody else) at home?

Similarly, the word değil not takes the predicative suffixes, as in Evde değilsin You are not at home.

The conjunction ve and is from Arabic and can be replaced with native-Turkish words like the postposition ile with and the particle de too.

Example. For Ayşe and David, we can write Ayşe ve David; Ayşe ile

David; and Ayşe, David de.

The ile can be suffixed, as in Ayşeyle.

1.7 Verbs

The dictionary-form of a verb is usually the infinitive, a verbal noun that can be analyzed as

stem + m@k;

otherwise, the stem itself is the dictionary head-word. Another verbal noun is stem + m@; I propose to call this a gerund, by analogy with the English gerund. (Lewis reserves ‘gerund’ for verbal adverbs.)

The Turkish infinitive does not have all features of the noun; for example, it does not take possessive suffixes. It does take some of the case-endings though. In particular, it forms a locative, to which a predicative suffix can be added to make a finite verb:

Example. Roman yazmaktayım I am engaged in writing a novel.

The gerund does have all of the features of nouns (except perhaps that it does not have a plural). In particular, it takes possessive suffixes:

Example. Gitmem l^azım My going is necessary, that is, I have to go.

From the gerund can be formed an adjective in -l#, to which a predicative suffix can be added, yielding a finite, necessitative verb:

Example. Kalkmalısın You must get up; Gitmeliyiz We must go.

Other flavors of verb are formed on this pattern and will be discussed below.

All verbs are regular, except the defective verb with stem i-. This verb does not have all of the inflexions that regular verbs do; the inflexions that it does have are used with regular verbs to form compounds. Where there is no inflexion of i-, but the corresponding meaning is needed, then the stem ol- supplies the inflexion. Also, ol- in some inflexions is used like var in an expression of possession. But olmak can also be treated as a regular verb meaning to become.

I mentioned that there were five series of personal endings. Here they are, along with the personal pronouns:

     0  ben    sen      o      biz        siz     onlar
     1 -(#)m  -(#)n  -(s)i  -(#)m#z    -(#)n#z     -l@r#
     2 -(y)#m   -s#n (-d#r) -(y)#z     -s#n#z  -(d#r)l@r
     3  -m     -n  —  -k       -n#z      -l@r
     4 -@y#m  -@s#n     -@    -@l#m    -@s#n#z     -@l@r
     5  — -s#n           -#n, -#n#z   -s#nl@r

I propose the following names for these series (two of which I gave earlier):

1. possessive;

2. predicative;

3. personal (the term is ambiguous; the point is that the endings on this line do nothing but assign person and number);

4. optative;

5. imperative.

The optative and imperative endings might be called modal. They attach directly to verb-stems. The optative endings (also called subjunctive) are little used, except in the first person. Indeed, one might think that there was a single series of modal endings, indicating a wish or command:

               gideyim  let me go
               git  go

gitsin  let him go/may he go

               gidelim  let’s go
               gidin(iz)  you all go

gitsinler let them go/may they go

According to Lewis, in ancient times there were first-person imperative endings; the singular one influenced the optative, and the plural one became the optative. Some roles of the optative or subjunctive are taken by the conditional, described below.

Note that every regular stem can stand by itself as a second-person singular imperative verb. Also, the stem i- does not take a modal ending, so ol- is used instead:

Example. Başınız sağ olsun is a formula offered to somebody grieving over a death. The command Be quick! is given by Çabuk ol!.

As in the case of the necessitative verb, so generally, a stem must be lengthened to an appropriate base before a non-modal ending can be added; the base will express some distinctions of tense and aspect (as well as mood, in that one of the bases is conditional). Before looking at bases, I shall first look more at stems.

1.7.1 Stems

The two-fold distinction in English between affirmation and denial (between I did it and I didn’t do it) can in Turkish be seen as three-fold. By addition of the indicated suffixes, a stem can become:

• negative: -m@; or

• impotential: -@m@.

Any of the three stems so far can be rendered potential by suffixation of

with -@bil.

Example. The defectiveness (in particular, the lack of infinitives) of the English modal auxiliaries gives various possibilities for translation. The second column below gives English infinitives without to (as might follow the words I shall); the third column gives a finite form (as might follow I).

gelme-  not come  do not come
geleme-  not be able to come  cannot come
gelebil-  be able to come  can come
gelmiyebil-  be able not to come  may not come
gelemiyebil- possibly be unable to come may be unable to come

The potential form can be considered as a compound with the verb bil-, meaning know. Other compounds, less common, are possible; for example, compounds with dur- stop or gel- indicate continuous action. Such compounding might be said to indicate distinctions of aspect.

Moreover, simple negation is achieved, as noted, with -m@. The impotential suffix is the negation of the obsolete verb u- be powerful or able.

The Turkish verb does not make the distinction between permission and ability that English can make with may and can.

The dictionary-stem (the stem before addition of negative, impotential or potential suffixes) can sometimes be further analysed. If it is not just a verbal root, then it might be a noun or adjective with a verbal suffix, often

-l@.

Example. başla- begin (make a head) temizle- clean (make clean) köpekle- cringe (make like a dog)

After the root or the l@ in a dictionary-stem, extensions may be present that indicate distinctions of voice. (If we insist on this term, then Turkish has more voices than the active and passive of English; English does have other ways of conveying the additional distinctions of voice that appear in Turkish).

When extensions are present, they appear in the following order:

1. reflexive: -(#)n;

2. reciprocal: -(#)ş;

3. causative:

• -t after polysyllabic stems in -l, -r or a vowel, and

• -d#r in other cases, except:

• -#r, -@r or -#t occur after some monosyllabic stems, and

• some exceptional forms occur as well;

4. passive: -#l, after stems ending in a consonant other than -l (otherwise it has the same form as the reflexive).

The significance of an extension is not always obvious from its name; that’s why all of the different stems that feature them will generally appear in the dictionary.

Like the ancient-Greek middle voice, the reflexive extension may suggest a doing for and not just to oneself, and it may have the same effect (and even the same form) as a passive extension (although these two kinds of extensions can be used together as well).

Used together, the reciprocal and causative extensions make the repetitive extension -#şt#r (other suffixes are used to the same effect in a few cases).

Sometimes the same verbal root can be used transitively or intransitively: gez- can mean either tour (the countryside) or take a walk (in the countryside). Also, the causative extension can make an intransitive stem transitive: öl- means die, while öldür- means kill, and kayna- and kaynatmean boil: respectively, what water does in a pot on the stove, and what I do when I put the pot there.

Used with a transitive stem, the causative extension could be called factitive: öldürt- means have [somebody] killed. As here, more than one causative extension can be used.

Here are some more examples of distinctions of voice:

Example. We have bulaşık yıka- wash dishes, but yıkan- wash oneself or be washed; the latter meaning is also expressed by yıkanıl-. Also:

• buluş- meet and bulun- be present;

• araştır- investigate;

• uyan- awake, uyandır- awaken, uyandırıl- be awakened;

• Çiçekleri koparttırmayınız, literally, Don’t let anybody allow [their children] to pick the flowers.

1.7.2 Bases

Regular verbs have several participles, some of which are formed from a stem by addition of the following suffixes:

• present: -@n;

• future: -(y)@c@k;

• past: -mış;

• aorist (not used with negative stems):

– positive: -@r, -#r or -r;

– negative: -m@z.

The predicative endings can be added to all of these participles but the present, turning them into finite verbs of the corresponding tenses. Here, i- has only the past tense.

There is said to be an ancient verb yor- go or walk. It survives as a suffix -#yor, which can be attached to a stem; the result does not stand on its own as a participle, but can receive a predicative suffix, making a present tense.

The aorist tense denotes habitual action; the present, ongoing action. The one past tense given so far may be called inferential, as its use does not imply that the speaker witnessed the event described—unless -d#r is added.

Since the third-singular predicative ending is empty, the bases we have now are also complete sentences:

Okuyor  (S)he is reading
Okumakta (S)he is engaged in reading
Okuyacak  (S)he will read
Okur  (S)he reads
Okumaz  (S)he does not read
Okumuş  It seems that (s)he read
(Okumuştur  (S)he read)
Okumalı  (S)he must read

Two more bases are formed from stems by addition of the following characteristics:

• simple past: -d#;

• conditional: -s@.

These attach to the stems of all verbs, including i-. The resulting bases take the personal endings. We have then, for example:

Okudu [I saw that] (s)he did read
Okusa  If (s)he would read…

All verbs formed with predicative or personal endings might be said to have the indicative mood. In particular, saying that a conditional verb is indicative makes sense if one takes If A then B to mean just B or not-A.

All bases so far described are simple. On the stem i- are simple are also formed compound bases. The simple bases result from suffixing a characteristic -m#ş, -d# or -s@ only. The non-conditional simple bases can be rendered conditional by being compounded with that simple base. Thus the bases in i- are five:

                        imiş-  imişse
                        idi-   idiyse-

                        ise-

All verb-bases have now been described. The bases formed on i- must be completed with personal endings. Personal endings may be attached to regular bases, as noted. Alternatively, a complete verb formed on a simple base in i- can follow or be suffixed to a regular verb-base.

Example. By appending bases in i- to indicative stems in oku-, these forms can be produced:

okuyormuş    okuyordu    okuyorsa    okuyormuşsa    okuyorduysa
okuyacakmış  okuyacaktı  okuyacaksa  okuyacakmışsa  okuyacaktıysa
okurmuş      okurdu      okursa      okurmuşsa      okurduysa
okumazmış    okumazdı    okumazsa    okumazmışsa    okumazdıysa
okumuş imiş  okumuştu    okumuşsa    okumuş imişsa  okumuş idiyse
okumalıymış  okumalıydı
             okuduydu    okuduysa                   okudu idiyse
okusaymış    okusaydı

If it is not associated in this way with a regular verb-base, then a complete verb in i- must follow or be suffixed to a noun, pronoun, adjective or değil.

In a verb with two characteristics, the second being conditional, the personal ending may precede this.

The distinction between the present and the aorist characteristics is of aspect. The aorist indicates habitual action; the present, an action ongoing or envisaged. A third way to indicate a present tense is by appending type-I endings or a verb in i- to an infinitive in the locative case: the result indicates ongoing action only.

Example. From yaz- write we can form:

   • Romanlar yazarım, I write novels;

• Roman yazıyorum, I am writing a novel [now, or in the future];

• Roman yazmaktayım, I am engaged in writing a novel.

The inferential characteristic is originally a past-participial ending. Used in a regular base, it indicates past action. This action is inferred from present appearances, or reported by somebody else—unless the regular base is followed by -d#r or a verb in i-, in which case no suggestion of hearsay or inference is meant. In particular, the ending -m#şt#r has the same force as -d#. The form imiş- may not indicate past time at all, but only hearsay.

Example. Barış means peace; so:

• A newspaper headline could read Barış geldi;

• The story under that head might say Barış gelmiştir;

• Somebody who read the story might report Barış gelmiş, or possibly Barış gelmiş imiş, Peace is said to have come.

The pluperfect tense is usually formed with -m#ş (+) idi-, more rarely

with -d# (+) idi.

The interrogative particle is m#; it always starts a new word. It makes a sentence into a question with a evet/hayır (yes/no) answer, and it follows the word that could also be an answer to the question. If this word is the verb, then the base is what is followed by m#; the personal ending, or the form of i- in use, is attached to this.

Exceptions include the following. If the personal ending is -l@r, then m# follows this, unless -d#r is also used. The characteristic d# of the simple past is inseparable from an attached personal ending, so if there is one, then m# must follow this.

2 Glossary

Turkish words here ending with a hyphen are verbs. (Some dictionaries replace the hyphen with m@k, giving the infinitive.)

altın gold
ara- seek

at horse baba father baş head

beyaz white
bil- know

boş empty

bul- find
buz ice
çabuk quick

çık- go out or up

çıplak naked
çiçek flower
çocuk child

dağ mountain dur stop ev house

gel- come

gece night gez- tour, walk

git- go
gün day

hız speed

kalk- get up

kop- break off

koyun sheep
kayna- boil
köpek dog

küpe earring l^azım necessary oda room

oku- read
öl- die

sağ alive, right süt milk

şeker sugar
şiir poem
tanrı god
temiz clean

uyu- sleep yaz- write

yıka- wash
zaman time

References

[1] G. L. Lewis, Turkish Grammar, Oxford, 1967.

[2] Mehmet Hengirmen, Turkish Grammar For Foreign Students, Ankara: Engin, 2000.